# LLM4ED: Large Language Models for Automatic Equation Discovery

### v1

@@ 5 Conclusion
We introduce a novel equation discovery framework guided by LLMs. It is aimed at facilitating equation discovery across diverse domains, transcending the confines of specialist communities, and making LLM-guided discovery accessible to a broader range of users. The framework leverages the generation and reasoning capabilities of large models to automatically complete the generation and optimization of equations. We employ natural language-based prompts to guide LLM in conducting iterative optimization of self-improvement and genetic algorithms. The results indicate that the two exhibit a strong synergistic effect, effectively striking a balance between exploration and exploitation. In the one-dimensional ODEs experiments, our framework achieved nearly equivalent performance to the state-of-the-art and compared the impact of large model capabilities on the performance of mined equations. In the future, we can further improve the existing framework in two aspects. One aspect is to further explore the design of natural language-based prompts, combining prior knowledge to better narrow down the search space and improve search efficiency and equation mining accuracy. On the other hand, further combining more effective evaluation methods to solve more complex scenarios, for example, when observations are sparse and noisy.
@@ Appendix A Prompts
The prompts utilized for initialization, self-improvement, and GA are shown in Fig. 10, Fig. 11, and Fig. 12, respectively.
@@
Figure 11: Prompt utilized in self-improvement.
Figure 12: Prompt utilized in GA.

### v2

@@ 5 Conclusion
We introduce a novel equation discovery framework guided by LLMs. It aims to facilitate equation discovery across diverse domains, transcending the confines of specialist communities, and making LLM-guided discovery accessible to a broader range of users. The framework leverages the generation and reasoning capabilities of LLMs to automatically generate and optimize equations. We employ natural language-based prompts to guide LLMs in conducting iterative optimization using self-improvement and genetic algorithms. The results indicate that the two strategies exhibit a strong synergistic effect, effectively balancing exploration and exploitation. Experiments demonstrate that the proposed framework can discover effective and accurate equations (including both PDEs and ODEs), directly from data. More importantly, our framework achieved performance comparable to the SR benchmarks, and the discovered equations tend to be more physically reasonable and possess better generalization ability. The impact of LLM capabilities on the performance of mined equations is also thoroughly discussed. The proposed framework demonstrates great potential in applying LLMs to scientific discovery and providing interpretability for complex physical phenomena.
The current framework has the potential for enhancement in two aspects. First, the training corpus of LLMs encompasses descriptions of diverse intricate physical processes and phenomena; therefore, providing a depiction of the nonlinear system to be mined and information on physical variables could be crucial for LLMs to generate rational equation forms. Further exploration is required to design natural language-based prompts that incorporate prior knowledge to effectively reduce the search space and enhance search efficiency and equation mining precision. Second, additional experimental validation is necessary to determine the robustness of equation mining using LLMs, particularly in scenarios with sparse and noisy observations. Integrating more efficient evaluation techniques might be essential for addressing more intricate situations.
Acknowledgments
This work was supported and partially funded by the National Center for Applied Mathematics Shenzhen (NCAMS), the Shenzhen Key Laboratory of Natural Gas Hydrates (Grant No. ZDSYS20200421111201738), the SUSTech – Qingdao New Energy Technology Research Institute, the China Meteorological Administration Climate Change Special Program (CMA-CCSP) (Grant No. QBZ202316), and the National Natural Science Foundation of China (Grant No. 62106116).
@@ Appendix A Prompts
The prompts utilized for initialization, self-improvement, and GA are shown in Fig. 10, Fig. 11, and Fig. 12, respectively. Different colors represent different components in the prompt. Specifically, red represents task descriptions, blue represents instructions, green represents historical samples, and brown represents other hints and constraints.
@@
Figure 11: Prompt utilized for self-improvement.
Figure 12: Prompt utilized for evolutionary search.
@@ Appendix B ODE datasets
The ODE datasets are selected from ODEBench [41] with different skeletons. The information for each ode is shown in Table 6.